A vehicle originally costed $23,000, had an estimated useful life of 8 years (?), and an estimated salvage value of $3,000. After 4 years of straight-line depreciation the assets total estimated life was revised from 8 years to 6 years. There was no change in salvage value. What is the depreciation expense in year 5?  
Answer #1

Annual depreciation = (23000-3000)/8 = $2500 per year

Net book value at the end of 4th year = 23000 - 2500*4 = $13000

Depreciation expense for year = [ (13000-3000)/2 = $5000